The Rogers Corporation has a gross profit of $709,000 and $254,000 in depreciation expense. The Evans Corporation also has $709,000 in gross profit, with $46,200 in depreciation expense. Selling and administrative expense is $187,000 for each company.  

a. Given that the tax rate is 40 percent, compute the cash flow for both companies.
  

  

b. Calculate the difference in cash flow between the two firms.
Answer #1

a)

The Rogers Corporation:

Cash flow = (Gross profit - administrative expense - depreciation)(1 - tax) + depreciation

Cash flow = (709,000 - 187,000 - 254,000 )(1 - 0.4) + 254,000

Cash flow = 160,800 + 254,000

Cash flow = $414,800

Evans Corporation:

Cash flow = (Gross profit - administrative expense - depreciation)(1 - tax) + depreciation

Cash flow = (709,000 - 187,000 - 46,200 )(1 - 0.4) + 46,200  

Cash flow = 285,480 + 46,200  

Cash flow = $331,680

b)

Difference = 414,800 - 331,680

Difference = $83,120